OpenStudy (anonymous):

How do I solve this? 1 + (1 / y) = (6 / y^2)

OpenStudy (anonymous):

Do you mean this: \[1 +\frac{ 1 }{ y } = \frac{ 6 }{ y^{2} }\]

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

First take LCM. of Left hand side. like. \[1 + \frac{ 1 }{ 3 } = \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

multiply with y^2 what do you get?

OpenStudy (anonymous):

hmm... (y^3 / y^3) + (y^2 / y^3) = (6y / y^3) 1 + y^2 = 6y y^2 - 6y +1 = 0 stuck again...

OpenStudy (anonymous):

wtf???? well i would get y^2+y=6 -->\[y^2+y-6=0\]

OpenStudy (anonymous):

any ideas how to solve that now?

OpenStudy (anonymous):

Hmm, how'd you get "y2+y−6=0"? Doesn't look right to me (?)

OpenStudy (anonymous):

your equation was: \[1+\frac{1}{y}=\frac{6}{y^2}\] right? now if you multiplay y^2 you get the equation that i posted.

OpenStudy (anonymous):

\[1 +\frac{ 1 }{ y } = \frac{ 6 }{ y^{2} }\] \[\frac{ y+1 }{ y }=\frac{ 6 }{ y^2 }\] multiplying both the equation by y we get \[y+1= \frac{ 6 }{ y }\] again multiplying by y we get \[y^2+y=6\] \[y^2+y-6=0\] \[\left( y+3 \right)\left( y-2 \right)\] therefore y=-3 or y=2

OpenStudy (anonymous):

if you multiply by y2 you get, \[y^{2} + \frac{ y^{2} }{ y } = \frac{ 6 }{ y^{2} } \times y^{2}\] \[ y^{2} + y - 6 = 0\]

OpenStudy (anonymous):

Allright, finally got it! Thanks!

OpenStudy (anonymous):

\[y^{2} + 3y - 2y - 6 = 0\]\[(y+3) (y-2) = 0\]y = -3 or 2.

OpenStudy (anonymous):

Hmm, I solved y2+y−6=0 using the formula \[x = -(p/2) +- \sqrt((p/2)^2-q)\] How did you get: y2+3y−2y−6=0 ?

OpenStudy (anonymous):

he looked for a way to part the function in linear factors. 3y-2y=y so it is your original function

OpenStudy (anonymous):

@math0101 dont you know the factorization method? the que you asked is from factorization method ;)

OpenStudy (anonymous):

@math0101 Its a method to factorize.

OpenStudy (anonymous):

I get this calculation: y2+3y−2y−6=0 (y+3)(y−2)=0 What I don't get is where you got y2+3y−2y−6=0 from

OpenStudy (anonymous):

Ie, what's the step y2+y−6=0 y2+3y−2y−6=0 ?

OpenStudy (anonymous):

okay here you will get the idea http://www.purplemath.com/modules/factquad.htm

OpenStudy (anonymous):

thank you!